OpenStudy (anonymous):

Q-9 Find the solution set of x + y = 5, 2 x - y = 7 a. (1,4) b. (2,3) c. (4,1) d. (1,4)

OpenStudy (anonymous):

x+y=5...1 2x-y=7...2 from ...1 we can say x=5-y...3 subst...3 in ...2 2(5-y)-y=7 10-2y-y=7 10-3y=7 -3y=-10+7 y=-10+7/-3 y=1 thn subst y=1 ...3 x=5-1 =4 (4,1)